NaF reacts with Th(NO3)4 according to the following unbalanced chemical equation:

NaF + Th(NO3)4 → NaNO3 + ThF4

Calculate the mass in grams of the excess reagent remaining after the complete reaction of 1.26 g of NaF with 7.80 g of Th(NO3)4.
Answer #1

Molar weight of NaF is 42 g

Molar weight of Th(NO3)4 is 480.1 g

the balanced equation is

4NaF + Th(NO3)4 = 4NaNO3 + ThF4

Hence the ratio of NaF: Th(NO3)4 is 4:1 = 1:0.25

i.e. 1 mole of NaF will react with 0.25 mole of Th(NO3)4

1.26 g of NaF = 0.03 moles

7.8 g of Th(NO3)4 = 0. 016 moles

Hence, the given mole ratio of NaF: Th(NO3)4 = 0.03:0.016 = 1:0.533

0.03 moles of NaF wil react with  0.0075 moles (0.03 x 0.25 moles) of Th(NO3)4.

i.e. 0.016-0.0075 =  0.0085 moles of Th(NO3)4 is excess

So, the excess reagent is Th(NO3)4. To complete the reaction 0.0075 moles of Th(NO3)4 required. i.e. 3.6 g (0.0075 x 480.1 g) of Th(NO3)4 will react with 1.26 g of NaF.

The remaining amount of Th(NO3)4 = 7.6-3.6 = 4.8 g